5. Calculate the molality of C2H5OH in a water solution that is prepared by mixing 50.0 mL of C2H5OH with 108.4 mL of H2O at 20
Answer #1

Molality:

Molality is defined as the number of moles of solute dissolved in 1 kg of solvent. The molality can be expressed as follows:

                            media%2F9c7%2F9c749d34-d3f9-4443-8180-95

---------------------------------------------------------------------------------------------------------------------------------------------------

First, we find the mass of C2H5OH and mass of H2O, using its mass conversion factor as:

                             media%2F534%2F534fb253-0c24-49c9-975d-f9

                     

From the above the data, we can find the number moles of C2H5OH and mass of C2H5OH is 39.45 g, using its molar mass as the conversion factor as:

               media%2Fa47%2Fa47c2b1e-168c-4a5d-a6b0-a1

Therefore, we calculate the molality as:

                            media%2Fc46%2Fc461090a-4d96-4298-9347-8a

Therefore, the molality is 7.90m and the correct option is (d) 7.90 m
Answer #2

Mass of C2H5OH = volume x concentration = 50.0 x 0.789 = 39.45 g Moles of C2H5OH = mass /molar mass of C2H5OH = 39.45/46.07 =
Answer #3

The answer is: d. 7.90 m Mass of C2H5OH = volume x density of C2H5OH -50.0 x 0.789 - 39.45 g Moles of C2H5OH mass/molar mass
Answer #4

ms= d x v 

=0.789 x 50=39.4 g


Mwt ( C2H5OH) = ( 12 X2 ) + ( 6X1 ) + ( 16X1 ) = 46 g/mol


n =ms/Mwt 

= 39.4 / 46 = 0.8 mol

 since 1ml = 1 g so mass of H2O is 100/1000 =0.1Kg

molality = 0.8 / 0.1 = 8 m

answered by: T.mai